### v1

# Understanding Biases in ChatGPT-based Recommender Systems: Provider Fairness, Temporal Stability, and Recency

Abstract.
This study explores the nuanced capabilities and inherent biases of Recommender Systems using Large Language Models (RecLLMs), with a focus on ChatGPT-based systems. It studies into the contrasting behaviors of generative models and traditional collaborative filtering models in movie recommendations. The research primarily investigates prompt design strategies and their impact on various aspects of recommendation quality, including accuracy, provider fairness, diversity, stability, genre dominance, and temporal freshness (recency).
Our experimental analysis reveals that the introduction of specific system roles and prompt strategies in RecLLMs significantly influences their performance. For instance, role-based prompts enhance fairness and diversity in recommendations, mitigating popularity bias. We find that while GPT-based models do not always match the performance of CF baselines, they exhibit a unique tendency to recommend newer and more diverse movie genres. Notably, GPT-based models tend to recommend more recent films, particularly those released post-2000, and show a preference for genres like ‘Drama’ and ‘Comedy,’ and ‘Romance’ (compared to CF ‘Action’, ‘Adventure’) presumably due to the RecLLMs’ training on varied data sets, which allows them to capture recent trends and discussions more effectively than CF models.
Interestingly, our results demonstrate that the Simple and Chain of Thought (COT) paradigms yield the highest accuracy. These findings imply the potential of combining these strategies with scenarios that favor more recent content, thereby offering a more balanced and up-to-date recommendation experience. This study contributes significantly to the understanding of emerging RecLLMs, particularly in the context of harms and biases within these systems. It highlights the complexities and potential of prompt design in impacting recommendation outcomes and paves the way for future research in optimizing accuracy, fairness, and user experience in such systems.
Recommender Systems, Large Language Models, Bias and Fairness in RS, Movie Recommendation Analysis, Prompt Design Strategies, ChatGPT, Stability and Diversity in Recommendations
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXX
1. Introduction
Recommender systems have increasingly become integral in various domains, offering substantial benefits for stakeholders in two-sided, or broadly, multi-stakeholder markets (Deldjoo et al., 2023; Abdollahpouri and Burke, 2021; Ekstrand et al., 2019; Zheng, 2019). These systems alleviate user information overload and enhance producer sales and revenue. From a learning viewpoint, recommendation models can generally be classified into two categories: ‘discriminative’ and ‘generative’ models (Deldjoo et al., 2022).
Discriminative models, often considered the traditional approach in recommendation systems, function by categorizing and differentiating to inform decisions about user preferences. For example, in the movie domain, these models can be compared to expert critics that are adept at discerning the distinctive characteristics that separate one type of item (movie) from another, given the audience. In other words, this differentiation is informed by factors such as user historical ratings, genre preferences, and other relevant attribute of items. Mainstream collaborative filtering (CF) recommendation models, such as matrix factorization (MF)-based models and more complex neural variations such as NeuMF (He et al., 2017), operate on this principle. Similarly, graph-based models, such as NGCF (Wang et al., 2019) and LightGCN (He et al., 2020), leverage the relational structure of the data to improve the quality of recommendations. Regardless of the methodological approach, traditional discriminative models are heavily based on knowledge derived from limited datasets and specific domains to calculate recommendations and inform decisions.
Generative models represent a transformative approach, focusing on understanding the underlying structure and distribution of data to generate authentic-looking content. Their application in recommender systems (RS) exemplifies a shift from simple preference matching to a more nuanced understanding of user behavior and interests. In the context of movie recommender systems, these models can be likened to hypothetical curators who not only distinguish between various items but also comprehend the overarching patterns and tastes of the audience. Thus, these models go beyond mere differentiation; they synthesize and anticipate user preferences, providing tailored recommendations.
Large Language Models (LLMs), exemplified by the Generative Pretrained Transformers (GPT), recently heralded a new era in Natural Language Processing (NLP). These Transformer-based models, such as BERT, GPT series, PaLM series, etc., are characterized by their vast number of parameters and are trained on extensive datasets using self/semi-supervised learning techniques. Their training on a wide range of internet text data endows LLMs with the exceptional ability to generate contextually relevant and coherent text in response to input prompts. Different from traditional recommendation systems, LLM-based models excel in capturing contextual information, comprehending user queries, item descriptions, and other textual data more effectively (Geng et al., 2022). Consequently, specifically-designed LLMs for recommender systems (RecLLMs) can harness LLMs to extract high-quality representations of textual features and leverage extensive external knowledge encoded within LLMs (Liu et al., 2023) to offer more nuanced insights into understanding user preference and profile. This is a significant leap forward compared to traditional RS. This aligns well with the overarching goal of delivering personalized, context-aware recommendations that reflect individual tastes and behaviors.
In general, LLMs can contribute to recommender systems in three principal ways (Wu et al., 2023; Li et al., 2023b; Xu et al., 2024). First, LLMs serve as direct recommenders by generating personalized content and suggestions, stemming from their advanced understanding of context and user queries (Geng et al., 2022; Wu et al., 2023). Second, they enhance traditional models by extracting rich semantic representations from textual data. This aspect is particularly important in addressing the data sparsity issue in historical interactions (Da’u and Salim, 2020). Third, LLMs act as complex simulators in the recommendation process, effectively testing and refining the system’s predictions (Liu et al., 2023). The current work in hand primarily concentrates on the first application, that is using LLMs as recommender, with ChatGPT serving as the case study.
The primary aim of this research is to enhance our understanding of “Provider Fairness within ChatGPT-based RecLLMs”. We focus on uncovering and understanding a range of biases – mostly underexplored in existing literature –, and examining how these biases manifest in comparison to traditional Collaborative Filtering (CF) models. This study looks into the dynamics of these biases, providing a detailed contrast between the two types of systems, and thereby contributing to a more nuanced understanding of the complexities inherent in modern RS.
In particular, this research diligently identifies and analyzes inherent biases inherent in these two types of systems, such as (i) the preference for popular movies, (ii) the balance between old and recent movie recommendations, (iii) the tendency to recommend certain movie genres, (iv) the consistency/stability of recommendations over time. For instance, GPT-based models were found to have a penchant for suggesting recent films like those released post-2000, exemplified by their frequent recommendations of newer titles such as ”Inception” (2010) and ”The Dark Knight” (2008), compared to CF models that often favored classics like ”The Godfather” (1972) and ”Casablanca” (1942). Furthermore, genre dominance was evident, with GPT models showing a distinct inclination towards ‘Drama’ and ‘Comedy’, whereas CF models predominantly recommended ‘Action’ and ‘Sci-Fi’ films. Another crucial aspect of this research is exploring the impact of prompt engineering in a zero-shot learning setting on the performance of RecLLMs in terms of personalization, diversity, recency, and fairness. Zero-shot learning refers to the model’s capacity to handle tasks for which it has not been explicitly trained, relying instead on its general language understanding and adaptability.
Here our study seeks to provide an answer to the following question: Can prompts serve as a controllability tool to steer the recommendation process in a specific direction (increase accuracy, reduce provider unfairness) based on the interests of providers or other stakeholders, and what potential harms are associated with this approach?
Figure 1. Conceptual idea behind our work
Example. We illustrate the impact of different “prompt design” strategies on the output of ChatGPT-based recommendation systems through a clear example, as shown in Figure 1. We use a randomly chosen user profile from the MovieLens dataset to demonstrate this effect. The user’s historical interactions with the system include a diverse range of movies such as “Toy Story” (categorized with genres Adventure—Animation—Children—Comedy—Fantasy), “If Lucy Fell” (Comedy—Romance), and “Hard Target” (Action—Adventure—Crime—Thriller).
To highlight the differences in recommendations based on the design of the system, we designed a total of seven prompts categorized into three distinct classes. These prompts were tailored to guide the recommendation system in different directions, showcasing how even with the same underlying user data, the system’s outputs can vary significantly depending on the design strategy employed.
  1. (1)
[…]
  2. (2)
    Beyond-Accuracy-Oriented Strategy. The goal of this beyond-accuracy-oriented approach is to broaden the user’s viewing experience by introducing diversity and novelty. It steers away from strictly aligning with known preferences and instead presents a variety of unique and perhaps surprising movie recommendations. It could be noted that for the same given query, this results in the user being recommended a movie like “Moon (2009)” - (genres: Drama—Mystery—Sci-Fi), which may be outside the user’s typical viewing history but offers a new cinematic perspective.
  3. (3)
    Reasoning-Oriented Strategy. This strategy is centered on enhancing user engagement and understanding. It not only recommends movies but also provides detailed explanation and reasoning regarding the suggestions. While using these explanations can potentially assist users in exploring and understanding why certain movies are recommended, we are particularly investigating this scenario since previous research has shown that motivating the LLM to reason, for example, through the use of a chain-of-thought (COT) reasoning, could result in more accurate responses from LLMs (i.e., relevant recommendations). An example of recommendation under this strategy could be “The Shawshank Redemption (1994)”, accompanied by an explanation of its themes of hope and redemption that resonate with the motivational depth found in “Toy Story.”
Promoting fairness in RecLLM: To combat popularity bias, we tested two approaches: (1) Instructional prompts: Statements like “Focus on fair recommendations, balancing popular and lesser-known movies” were integrated into various scenarios, aiming to influence RecLLM’s output. (2) Role-based directives: We explored directly incorporating fairness into the “system’s role”, testing its effectiveness compared to solely instructional prompts. This experiment seeks to identify the most effective strategies for achieving fair recommendations in RecLLM. In sum, our work demonstrates how strategically designing a prompt’s semantic intent and structure can impact a recommender system’s performance across accuracy, provider fairness, diversity, and temporal freshness (recency).
Contributions. The following is a list of proposed contributions in the current work:
  1. (1)
    Enhanced Analysis of Prompt Design in Zero-Shot RecLMMs. This research provides an in-depth study into how the design of prompts (in terms of intent, structure, and system role) affects the performance of ChatGPT-based recommendation systems. It investigates how different prompt design impacts factors such as accuracy, fairness, diversity, and temporal freshness in recommendations. This study involves experimental designs that compare various prompt strategies (e.g., accuracy-oriented, beyond-accuracy-oriented, explanation-oriented) and analyze their impacts on the recommendation of niche versus popular content, as well as the balance between old and recent movies.
  2. (2)
    Identification and Analysis of Different Biases in Zero-Shot RecLLMs. This work identifies and analyzes numerous biases that may be present in RecLLMs such as ChatGPT, thus laying the groundwork for investigating bias mitigation strategies. Key biases identified under scrutiny include popularity bias, genre preference bias, and temporal (recency) bias. We study how these biases uniquely manifest in LLM-based systems versus traditional Collaborative Filtering (CF) models.
  3. (3)
    Stability Analysis in Zero-Shot RecLLMs: This world also studies the stability of recommendations provided by LLM-based systems over time. Stability in this context refers to the consistency of recommendations given similar input conditions at different times. Our study analyzes and discusses factors that might impact stability (besides the parameters controlled by LLMs), such as changes in underlying data, model updates, and prompt variations. It would also explore the implications of stability for user satisfaction and trust in the system in real-life deployments.
The structure of the paper is organized as follows. Section 2 presents the related work, focusing on the exploration of fairness in recommendation systems (RS) and pre-trained language models (LMs), as well as their applications in enhancing RS. Section 3 is dedicated to presenting a thorough evaluation of ChatGPT. This is this research’s core contribution which we present through the suite of goal-oriented prompts’ (Section 3.1), the repeated experiments for stability analysis (Section 3.2), an understanding of the system’s role in ChatGPT-based RecLLM (Section 3.3), and the promotion of fairness (Section 3.4). Additionally, we explore the explicit versus implicit scenario, which is indicated in prompt structuring (Section 3.5). Following this, Section 4 outlines the experimental setup employed in our study. The results and key findings of these experiments are thoroughly discussed in Section 5. The paper concludes with Section 6, where we summarize our findings and outline potential avenues for future research.
2. Related work
In this section, we briefly review some related work on recommendation systems and LLM techniques.
2.1. Fairness in Recommender Systems
Fairness has emerged as a pivotal topic within the AI community in the last few years, arguably becoming the most scrutinized aspect across various branches of trustworthy AI, including fairness, security, privacy, and explainability. In RS, fairness has garnered significant attention due to the multi-stakeholder nature of these systems (Deldjoo et al., 2023; Ekstrand et al., 2019). Unfairness, even in its minimal form, can adversely impact various stakeholders, including consumers, producers, system designers, supply chains, and even the environment – the latter are often referred to as ‘side-stakeholder’.
The body of literature on fairness in RS is diverse, encompassing multiple perspectives and dimensions. Recent surveys (Ekstrand et al., 2019; Deldjoo et al., 2023; Amigó et al., 2023), have introduced various taxonomies to categorize these diverse aspects into several orthogonal or partially-orthogonal dimensions. Key dimensions recognized in the literature include the main stakeholder in question (e.g., consumer vs. producer), the target benefits associated with each (such as effectiveness vs. item exposure), the granularity of sensitive groups for assessing fairness (individual vs. group level), and other dimensions including the core definition of fairness, temporal aspects, and more. These dimensions offer a comprehensive view of the fairness landscape in RS, as detailed in these surveys.
In the context of the present study, and with regard to positioning this work within the literature of fair Recommender Systems (fair-RS), we have introduced Table 1. This table is designed to categorize existing literature along two principal dimensions: the stakeholder in question (consumer vs. producer) and the nature of the core RS under scrutiny (traditional RS vs. those based on recent advancements in LLMs). We briefly review these dimensions in the following.
Core RS under scrutiny. We distinguish between ‘traditional’ RS and those enhanced by ‘RecLLM’. This distinction is crucial because, while RecLLM, such as those based on GPT-like architectures, promise to significantly advance RS landscape with more nuanced and personalized recommendations, they also raise concerns about inherent biases. The vast and unregulated nature of internet data used for training LLMs raises concerns about inherent biases against specific races, genders, popular brands, and other sensitive attributes. For example, if an LLM is predominantly trained on data from popular e-commerce sites, it might disproportionately recommend products from more recognized brands, overlooking niche or emerging brands. Similarly, biases in language around gender or race could skew recommendations in subtle but impactful ways. Therefore, accurately measuring these biases is a crucial initial step, and forms the key goal of our current work, in developing effective mitigation strategies, ensuring that advancements in RecLLM do not amplify existing inequalities. This is the key goal of the current work in hand.
  • Traditional RS. This column lists studies that have focused on traditional methods of recommender systems. They primarily rely on CF algorithms (possibly using side information of users and items) without the advanced natural language processing capabilities of LLMs. Within these works, some are dedicated to building evaluation frameworks for evaluating RS unfairness, while others focus on developing various mitigation strategies.
  • RecLLM: This column spotlights the burgeoning research domain that merges language models (LMs) with RS representing a significant shift towards harnessing advanced NLP techniques to enhance the accuracy and relevance of recommendations. These studies explore the use of various LMs, including BERT-based models (Shen et al., 2023) and recent LLMs such as GPT-like architectures (Li et al., 2023a; Zhang et al., 2023). Additionally, beyond the scale of LMs, the target tasks within RS–such as classical recommendation (top-$k$ ranking (Li et al., 2023a; Zhang et al., 2023), sequential), conversational RS (Shen et al., 2023), explanation generation, multi-modal recommendations–provide dimensions that could be used for further categorizing these works.
Table 1. Research landscape in Recommender Systems focusing on Consumer and Producer fairness with respect to Age, Gender, Other, and Others
Traditional RS RecLLM
Consumer Fairness Activity (Li et al., 2021; Hao et al., 2021; Xiao et al., 2020; Rahmani et al., 2022b)
Demographics (Deldjoo et al., 2021a, b; Wu et al., 2021b; Weydemann et al., 2019; Farnadi et al., 2018) (Shen et al., 2023; Zhang et al., 2023)
Merits (Sühr et al., 2021; Gómez et al., 2021)
Others (Wan et al., 2020; Lin et al., 2021)
Producer Fairness Popularity (Dong et al., 2021; da Silva et al., 2021; Ge et al., 2021; Zhu et al., 2021) (Li et al., 2023a)
Demographics (Kırnap et al., 2021; Boratto et al., 2021)
Price/Brand/Location (Deldjoo et al., 2021a; Burke et al., 2018; Liu et al., 2020; Shakespeare et al., 2020)
CP Fairness Mixed attributes (Naghiaei et al., 2022; Rahmani et al., 2022a; Chakraborty et al., 2017; Patro et al., 2020; Wu et al., 2021a; Do et al., 2021)
Stakeholder. As mentioned earlier, a dominant aspect that can be utilized to classify almost all literature on group fairness is the market side focus—whether they concentrate on a single-side market (defined either by consumers or providers) or on both sides. Within each of these segments, as you can observe, we can further categorize the literature based on which sensitive attribute groups are defined. For example, sensitive attributes such as the demographics of consumers (age, gender) and producers, as well as the popularity of items (on the produce side) are quite common focal points.
  • Consumer Fairness. This category is subdivided into various attributes like Activity, Demographics, Merits, and Others. It includes studies that focus on ensuring fairness among consumers of the recommender system based on these attributes. For instance, ensuring that recommendations are not biased towards a particular demographic group.
  • Producer Fairness. This focuses on the fairness towards the providers or producers of the content or products recommended by the system. It includes subcategories like Popularity, Demographics, and Price/Brand/Location. These studies might address issues like ensuring lesser-known or niche producers get fair visibility and opportunity in the recommendation process.
  • CP Fairness. This category involves studies that consider both consumer and producer fairness simultaneously, addressing the balance between the two.
It is noteworthy that while traditional Fair-RS research has been extensively explored, RecLLM is an emerging field, presenting its own unique considerations and challenges. Our work is situated in the ‘Producer Fairness’ category under ‘RecLLM’, focusing on producer fairness in the context of ChatGPT. It focuses on producer fairness in the context of ChatGPT, particularly examining how prompt engineering techniques can be leveraged to address or potentially enhance fairness.
The study by Zhang et al. (2023) evaluates the consumer fairness side of zero-shot GPT recommendations, focusing on a variety of consumer demographic attributes but not addressing producer-side fairness. Their work introduces a novel benchmark, FaiRLLM, for evaluating the fairness of RecLLM, highlighting ChatGPT’s biases towards certain sensitive user attributes in music and movie recommendations. Conversely, the work by Li et al. (2023a) aligns more closely with ours, focusing on producer unfairness however within the specific domain of news recommendation. This study investigates ChatGPT’s performance in news recommendation, exploring aspects like personalization, provider fairness, and fake news detection. Similarly, research such as that by Di Palma et al. (2023) examines the personalization and popularity bias in RecLLMs, particularly in the context of ChatGPT and other LLMs. However, these studies do not address the intricacies of prompt engineering in RecLLMs, nor do they comprehensively address the various forms of biases studies in the current study. Our research goes beyond examining provider fairness based on popularity bias and also considers other potential harms, such as the recency of recommended items and the stability of recommendations—aspects not explored in the aforementioned studies.
2.2. Leveraging Pre-trained LMs and Prompting for Recommender Systems
Recent advancements in recommender systems (RS) have been significantly influenced by the integration of LLMs and innovative prompting strategies. The use of natural language in recommendation tasks has been explored in various ways. For instance, Hou et al. (2022) utilize natural language descriptions and tags as inputs into LLMs to create user representations for more effective recommendations. This contrasts with the narrative-driven recommendations (Bogers and Koolen, 2017) that rely on verbose descriptions of specific contextual needs.
In terms of prompting techniques, early methods relied on few-shot prompting, where training examples are used as a guide for LLMs (Brown et al., 2020). With prompt learning, tasks are adapted to LLMs rather than the other way around, utilizing discrete prompts or continuous/soft prompts for task performance. This approach has shown promise across various tasks, including recommendation tasks.
Personalizing LLMs for recommendation is crucial for understanding a user’s intent and addressing their personalized needs. Recent efforts like P5 (Geng et al., 2022) and OpenP5 (Xu et al., 2023) have integrated several recommendation tasks into one LLM using personalized prompts. This approach reformulates recommendation tasks as sequence-to-sequence generation problems, demonstrating the flexibility of LLMs in handling diverse recommendation scenarios. Lastly, prompt transfer research, such as SPoT (Vu et al., 2021) and ATTEMPT (Asai et al., 2022), focuses on learning from source tasks and utilizing this knowledge for target tasks. This method, including knowledge distillation techniques, shows potential in intra-task prompt distillation and cross-task prompt transfer, contributing to the efficiency and effectiveness of LLM-based recommendation models.
3. Evaluation of ChatGPT-based RecLLM
Our work focuses on evaluating ChatGPT as a zero-shot recommender in the movie domain, aiming to extend beyond previous research by delving more deeply into biases and potential harms caused by such systems. Given the extensive nature of our experiments, we chose to concentrate on the single domain of movie recommendations. In this context, our evaluation methodology for ChatGPT was comprehensive and multifaceted. We conducted extensive experiments across a multi-dimensional testbed to assess the robustness and adaptability of the system. This evaluation particularly emphasized understanding the role and impact of ChatGPT in three key areas:
  1. (1)
    Goal-oriented prompt,
  2. (2)
    System role,
  3. (3)
    Promoting fairness, as well as
  4. (4)
    Implcit vs. explicit scenario
These dimensions were varied to observe their individual and combined effects on the quality of the recommendations provided. Please note that since the primary objective of this study is centered on (provider) fairness, the emphasis on fairness has been integrated either as a distinct scenario within the goal-oriented prompts or within the system role itself, and the impact is evaluated.
3.1. Goal-oriented prompts
Seven distinct prompt scenarios were designed, categorized into three classes focused on personalization, beyond-accuracy metrics, and reasoning. These classes aim to explore how variations in the nature of the prompt influence the recommendations. The scenarios are detailed in Table 2 and further exemplified in Section 1.
Table 2. Overview of Prompt Scenarios.
Scenario Description and Prompt
Personalization Focused
S1 - Simple
Basic recommendations without additional context.
Prompt: ”Recommend 10 movies that the user will likely enjoy.”
S2 - Genre-focused
Recommendations focusing on genres and themes similar to the user’s past favorites.
Prompt: ”Recommend 10 movies that the user will likely enjoy, particularly focusing on genres and themes similar to their past favorites.”
S3 - Rating-focused
Incorporation of the user’s explicit ratings into the recommendations.
Prompt: ”Recommend 10 movies the user will likely enjoy, taking into account both their favorite genres and past movie ratings.”
Beyond-Accuracy Focused
S4 - Diversify Recommendations
Suggesting lesser-known films to diversify the user’s experience.
Prompt: ”Suggest 10 high-quality, lesser-known films that diverge from mainstream blockbusters, yet align with the user’s tastes.”
S5 - Surprise
Offering unexpected recommendations for exploring new preferences.
Prompt: ”Offer 10 unique and unexpected movie recommendations aimed at broadening the user’s cinematic horizons beyond their usual preferences.”
Reasoning-Explanation Focused
S6 - Motivate Reasoning
Providing reasoning for each recommendation to enhance transparency.
Prompt: ”Provide 10 carefully selected movie recommendations, each accompanied by a rationale explaining its suitability for the user’s preferences.”
S7 - Chain of Thought (COT)
Engaging in a logical reasoning process to arrive at the recommendations.
Prompt: ”Let’s think this through: What would be 10 great movie recommendations for this user and why?”
In particular, scenarios S1 to S3 are focused on core personalization aspects, encompassing basic recommendations without additional context (S1), genre preferences (S2), and the incorporation of explicit user ratings (S3). Meanwhile, Scenarios (S4) and (S5) are designed to enhance the diversity and novelty of the recommendations, with S4 emphasizing lesser-known film recommendation and S5 aiming to broaden the user’s cinematic horizons with unique and unexpected choices. Scenarios S6 and S7 are explanation-motivation oriented, where S6 seeks to provide reasoning for each recommendation to enhance transparency, and S7 involves a logical ste-by-step reasoning process to arrive at the recommendations. These scenarios compose a competitive and insightful set of prompts, some of which such as COT, have been successfully tested in other ML disciplines (Wei et al., 2022; Wang et al., 2022; Chu et al., 2023).
Note. The CF baselines used in this work operate in an implicit setting. For consistency and fairness, all scenarios in the generative part adhered to this setting, thereby not revealing user movie ratings to the LLM in the prompts. Scenario ‘Rating-focused’ S3 is the only exception, included solely for completeness. This is discussed in Section 3.5.
3.2. Repeated Experiment for the Stability of the Analysis
The experiments were conducted at a temperature setting of 0.0, aiming for predictable responses. However, we observed variations in the recommendations, possibly due to factors like trending data, recent updates in the database, or changes in GPT’s behavior. This suggests that the recommendation system may interpret the same input differently in each iteration, independent of the temperature setting.
To assess the robustness and reliability of our system, we repeatedly conducted recommendation queries for each user, five times each, to examine if outputs vary despite consistent input.
We further adopted a “bootstrapping sampling strategy” during the evaluation stage, as detailed in Section 4.3. This method is particularly effective in revealing the reliability of the results, even when dealing with datasets that are small or may not perfectly represent the entire population. This thorough evaluation approach, therefore, enhances our confidence in the system’s performance across various potential real-world scenarios.
3.3. Understanding ChatGPT’s “System” Role
In this study, we sought to understand the impact of different system roles assigned to ChatGPT on recommendation outcomes. We focused on assessing whether responses significantly differ when specific “system” instructions are attached, and how these variations influence the functionality of a recommender system. Specifically, we aimed to determine the effectiveness of embedding fairness directly into the system’s operations versus incorporating it in the prompt structure. Table 3 summarizes the roles assigned to ChatGPT during our experiments:
Table 3. Summary of Different Roles Assigned to ChatGPT during experiments
Role ID Description
R0 - No Role
Direct user-driven prompts without system context.
R1 - System Role as Recommender
System-centric role focusing purely on user information.
R2 - System Role as Fair Recommender
System role with an explicit fairness objective.
The roles explored provide insight into how ChatGPT’s framing, in terms of its responsibilities and objectives, shapes its behavior in making recommendations. This analysis is crucial for comprehending the subtleties of human-AI interaction within diverse contexts and for distinct purposes, especially in how the system interprets and adheres to varying directives.
3.4. Promoting Fairness
We also aim to understand whether integrating a fairness statement in the system or in the prompt itself is more effective. To assess the impact of explicit directives on the recommendation system, each scenario was optionally combined with a ‘Fairness Emphasis Statement.’ These statements directed the model towards specific objectives, here ensuring fairness. The considered emphasis options included
  • E0 - Without Fairness Statement: No fairness statement is included in this option. Prompts are used as described in Table 2.
  • E1 - With Fairness Statement: This option emphasizes ensuring fair recommendation between popuar nd less popular movies. “Prompt: Ensure a fair representation of both popular and less-known movies. Based on these movies: {user_movies_string}, recommend 10 movies that the user will likely enjoy”. The latter prompt is combining Scenario (S1/E1).
3.5. Explicit vs. Implicit Scenario
This test evaluated whether displaying the movies a user has rated, as indicated in {user_movies_string}, enhances the recommender system’s quality, based on the premise that including such information could improve recommendations. Scenarios S1 to S6, except S3, were conducted implicitly, without user ratings in the prompts, mirroring the implicit mode of the CF baseline experiments. Uniquely, S3 was examined in both explicit and implicit contexts. In the explicit context, user ratings were integrated into the prompts, but were absent in the implicit context. This explicit approach was specific to S3. An example of {user_movies_string} and the presentation of user movie lists and ratings in the explicit context is as follows.
Example. To illustrate the experimental scenarios, consider a user’s movie list represented in this manner.
The Matrix (Genres: Action|Sci-Fi), Inception (Genres: Action|Thriller)
In the Explicit scenario, the prompt would be constructed to include these ratings, exemplified here:
The Matrix (Genres: Action/Sci-Fi, Rating: 5/5), Inception (Genres: Action/Thriller, Rating: 4.5/5)
The primary hypothesis under investigation is whether the inclusion of explicit ratings enables the system to tailor recommendation more closely to the user’s demonstrated preferences, thereby enhancing the personalization of the recommendation process.
4. Experimental Setup
This section outlines experiments conducted to assess various reproducibility aspects, including the obtained results and our observations.
4.1. Datasets and splitting
In this study, we utilized the MovieLens-Latest-Small dataset provided by the GroupLens group, [^1] https://grouplens.org/datasets/movielens/ which contains the most recent versions of movies. It is important to note that we intentionally focused our research within the domain of movies. This decision was made due to the potential varying capabilities of GPT-based models across different domains. Additionally, conducting extensive experiments with GPT models is time and resource-intensive, which is why we opted to work with this specific dataset. We acknowledge that there is potential for future exploration in other domains, and we leave this avenue open for future research. Detailed dataset parameters, including $\frac{R}{U}$ (ratings per user) and $\frac{R}{I}$ (ratings per item), resulting in a higher density $\frac{R}{U\times I}$, are presented in Table 4 for reference. Our setting involves a train-validation-test split for the data with ratios 80%, 10%, and 10%, respectively.
4.2. Evaluation Metrics
Our study employs a diverse range of metrics to assess the performance of our recommendation system. These metrics are categorized into two primary classes: Accuracy, Beyond-accuracy and Provider Fairness.
Accuracy Metrics. These metrics measure the quality of the top-$k$ ranking in the recommendation list.
  • NDCG@k. A measure of ranking quality, particularly the placement of relevant items.
  • Recall@k. The proportion of relevant items that are successfully recommended.
High scores in these metrics indicate more accurate and relevant recommendations.
Beyond-Accuracy and Provider Fairness Metrics. Beyond conventional accuracy metrics, these metrics evaluate the diversity, and fairness of the recommendations w.r.t provider perspective. They assess how well the system introduces new and varied content and how equitably it treats different item providers.
  • Gini Index: Measures the inequality in the distribution of item recommendations. A lower Gini Index indicates a more equitable distribution among items. The Gini Index is beneficial for identifying monopolization of recommendations by a few items, but it may not always capture the full picture of item diversity.
  • HHI (Herfindahl-Hirschman Index): Assesses the concentration of recommendations among items. A lower HHI value suggests a healthier, more diverse recommendation landscape. This index is particularly useful for detecting dominance by a small number of items, though it might overlook the distribution nuances among less recommended items.
  • Entropy: Reflects the diversity and unpredictability of recommendations. Higher entropy values indicate greater diversity. Entropy is advantageous for understanding the overall unpredictability of recommendations but may not directly indicate the fairness of item exposure.
Each of these metrics has its own advantages and limitations, making them collectively essential for a well-rounded evaluation. They help in understanding not just the effectiveness of the recommendation system in terms of user satisfaction (accuracy) but also its robustness in promoting diverse and fair content distribution.
4.3. Bootstrapping Sampling Strategy
In our evaluation stage, we employed a bootstrapping sampling strategy, a statistical technique where 1000 samples were generated by repeatedly resampling our dataset with replacement. This method was chosen to enhance the robustness and reliability of our analysis, particularly in measuring the performance metrics of our recommendation system, such as NDCG, Recall, and Precision.
Through this approach, we computed means for each metric across all bootstrap samples, then calculated their averages and 95% confidence intervals. This strategy is implemented for understanding the variability and confidence in our metrics, ensuring our evaluation reflects the system’s likely real-world performance. It is especially beneficial in situations with small or non-representative samples, allowing for more reliable population inferences.
4.4. CF Baselines.
The choice of models and the corresponding hyperparameter optimization process is outlined as follows.
Table 4. Statistics of the final datasets used in this work after $k$-core pre-processing.
Dataset —U— —I— —R— $\frac{R}{U}$ $\frac{R}{I}$ Density Item Gini User Gini
MovieLens Latest Small 610 9,724 100,836 165.30 10.36 98.3% 0.715 0.603
4.4.1. Models.
We use a suite of competitive, CF recommendation models as baseline ranking models in our post-processing approach, as summarized below.
  • BPR (Rendle et al., 2012): A conventional recommendation model that employs matrix factorization to learn user and item embeddings of low dimensionality, and optimize the model based on the pairwise ranking of items for each user to predict whether a user prefers a given item over another;
  • ItemKNN (Sarwar et al., 2000): An item-based K-nearest employs neighbors algorithm, utilizing similarity metrics such as cosine similarity to identify the closest item neighbors. These neighboring items are then leveraged to predict scores for user-item pairs.
  • MultiVAE (Liang et al., 2018): A non-linear probabilistic deep learning model that extends a variational autoencoder (VAE) structure to collaborative filtering for implicit feedback, and acquires the underlying representations of users and items from their interactions to create recommendations in an unsupervised way.
  • LightGCN (He et al., 2020): A pure collaborative filtering method that utilizes a simplified version of graph convolutional networks (GCNs) without nonlinear activation functions and additional weight matrices. It learns user and item embeddings through graph propagation rules and user-item interactions, making it scalable and efficient.
  • NGCF (Wang et al., 2019): A graph-based recommendation model that employs a neural network architecture and learns high-order connectivity and user-item signals based on the exploitation of the user-item graph structure, by propagating embeddings on it.
The selected CF models constitute a set of competitive baselines from various natures, representing a diverse array of recommendation approaches (classical, neural, graph-based). Additionally, the TopPop method is included as a non-informative baseline, as an essential baseline to provide a case where the outcomes are heavily biased towards popular items.
4.4.2. Hyperparameter Tuning.
In our research, the primary focus is on the hyperparameter tuning of various collaborative filtering (CF) recommendation models, selected as baselines. These models are specifically tailored and evaluated within implicit feedback scenarios, where the subtleties of user preferences are inferred from indirect interactions. This approach ensures a comprehensive exploration of hyperparameter spaces to optimize model performance in these nuanced environments. The RecBole public library [^2] https://www.recbole.io/ is used for implementing and applying these models. Hyperparameter tuning is an essential step in optimizing the performance of these models. We employ a greedy search strategy to identify the best configurations based on their performance on the validation set. The following is a summary of the hyperparameter tuning conducted for each model:
  • BPR-MF. Embedding size choices ’[32, 64, 128]’ and learning rate choices ’[1e-4, 5e-4, 1e-3, 5e-3]’. Total of 12 cases.
  • ItemKNN. k choices ’[10, 50, 100, 200, 250, 300, 400]’ and shrink choices ’[0.0, 0.1, 0.5, 1, 2]’. Total of 35 cases.
  • NGCF. Learning rate choices ’[1e-4, 5e-4, 1e-3]’, hidden size list choices ’["[64, 64, 64]", "[128, 128, 128]"]’, node dropout choices ’[0.0, 0.1, 0.2]’, message dropout choices ’[0.0, 0.1, 0.2]’, and regularization weight choices ’[1e-5, 1e-3]’. Total of 108 cases.
  • MultiVAE. Learning rate choices ’[1e-4, 5e-4, 1e-3, 5e-3]’ and latent dimension choices ’[64, 128, 200, 300, 512]’. Total of 20 cases.
  • LightGCN. Learning rate choices ’[5e-4, 1e-3, 2e-3]’, number of layers choices ’[1, 2, 3, 4]’, and regularization weight choices ’[1e-05, 1e-04, 1e-03, 1e-02]’. Total of 48 cases.
The total number of hyper-parameters is 14, encompassing a total of 223 experimental cases. This approach ensures that we effectively explore a wide range of hyperparameter configurations, particularly for models operating in implicit feedback scenarios, which are crucial for our resear
5. Results
In this section, we present insights gained from the experimental validation of our system. We have divided our analysis into four key areas, each focusing on different aspects of performance: (i) personalization (cf. Section 5.1), (iii) stability of recommendations over different runs (cf. Section 5.2), (ii) provider fairness and diversity of recommendations (cf. Section 5.3), and finally (iv) the temporal freshness (or recency) of recommendations (cf. Section 5.4). We also provide an interesting discussion on the genre dominance/bias of recommendation models employed in Section 5.5. We approach this section by formulating a number of the hypotheses for each part, afterward showcasing the results, and providing the associated analysis. Finally, we will provide answers to these hypotheses.
5.1. Personalization of Recommendations
This section focuses on the personalization performance measured by recommendation top-$K$ accuracy. Results can be found in Table 5, in which each generative scenario (i.e., using ChatGPT) could be found on the upper tabs while the result of CF baselines is on the bottom. It should be noted that, given the extensive nature of these experiments, not every possible scenario combination was tested independently; rather, a select subset of scenarios was examined.
We will analyze the results by evaluating the average performance (abbreviated as Avg Perf.) in terms of NDCG and Recall across different conditions/models. These conditions include exploring the impact of (i) the system’s role: No Role (R0), Standard Recommender (R1), and Fairness-aware Recommender (R2), and (ii) the presence (E1) or absence (E0) of a fairness statement in the LLM prompts.
Table 5. Table of the performance result measured in terms of recommendation Top-10 accuracy.
“System” Role: No (R0), Emphasis: No (E0)
Avg Perf. Bootstrap Mean Bootstrap Conf
Type Model NDCG Recall NDCG Recall NDCG Conf. Recall Conf.
Generative (ChatGPT) S1. Simple 0.008803 0.00994 0.008784 0.010011 (0.006323, 0.011484) (0.00674, 0.013824)
S2. Genre-focused 0.006964 0.007672 0.006994 0.007689 (0.004881, 0.009200) (0.005079, 0.0106819)
S4. Diversify 0.001906 0.002946 0.001896 0.002946 (0.000822, 0.003185) (0.001559, 0.004582)
S5. Surprise 0.007265 0.006968 0.007286 0.00696 (0.004671, 0.010279) (0.004130, 0.010021)
S6. Motivate Reasoning 0.010152 0.010906 0.010073 0.01098 (0.007446, 0.013124) (0.007511, 0.0148920)
S7. Chain-of-thought (COT) 0.006573 0.005848 0.006554 0.005821 (0.004513, 0.008940) (0.003853, 0.008015)
“System” Role: No (R0), Emphasis: Fair (E1)
Avg Perf. Avg. Bootstrap Bootstrap Cnf
Type Model NDCG Recall NDCG Recall NDCG Recall
Generative (ChatGPT) S1. Simple 0.010431 0.010557 0.010404 0.010579 (0.007491, 0.013792) (0.007295, 0.014063)
S2. Genre-focused 0.005221 0.005471 0.005211 0.00555 (0.003413, 0.007300) (0.003361, 0.008196)
S4. Diversify 0.001717 0.003029 0.001713 0.003024 (0.000607, 0.003261) (0.001504, 0.005083)
S5. Surprise 0.004095 0.005089 0.004059 0.005112 (0.002399, 0.006132) (0.002645, 0.008248)
S6. Motivate Reasoning 0.004315 0.005097 0.004302 0.005031 (0.002246, 0.006819) (0.002668, 0.007909)
S7. Chain-of-thought (COT) 0.007207 0.009344 0.007177 0.009373 (0.004493, 0.010162) (0.005808, 0.013662)
“System” Role: Normal Recommender (R1), Emphasis: Fair (E1)
Avg Perf. Avg. Bootstrap Bootstrap Conf
Type Model NDCG Recall NDCG Recall NDCG Conf. Recall Conf.
Generative (ChatGPT) S1. Simple 0.013007 0.012606 0.012984 0.012514 (0.009799, 0.016379) (0.008610, 0.016971)
S2. Genre-focused 0.00854 0.00987 0.008582 0.00983 (0.006277, 0.010984) (0.006564, 0.013434)
S4. Diversify 0.00314 0.005606 0.003162 0.005523 (0.001477, 0.005548) (0.003086, 0.008230)
S5. Surprise 0.006473 0.006435 0.006403 0.00642 (0.004065, 0.009319) (0.003627, 0.009950)
S6. Motivate Reasoning 0.009379 0.00957 0.009447 0.009496 (0.006160, 0.012722) (0.006056, 0.013061)
S7. Chain-of-thought (COT) 0.013571 0.012887 0.013622 0.012941 (0.009836, 0.018061) (0.009306, 0.017057)
“System” Role: Fair Recommender (R2), Emphasis: No (E0)
Avg Perf. Avg. Bootstrap Bootstrap Cnf
Type Model NDCG Recall NDCG Recall NDCG Conf. Recall Conf.
Generative (ChatGPT) S1. Simple 0.014268 0.015796 0.014257 0.015731 (0.011243, 0.017614) (0.011896, 0.019893)
S2. Genre-focused 0.008847 0.01132 0.00883 0.011382 (0.006700, 0.011233) (0.007998, 0.015372)
S4. Diversify 0.002844 0.004901 0.002825 0.004872 (0.001303, 0.004922) (0.003205, 0.006746)
S5. Surprise 0.007504 0.00729 0.007446 0.007304 (0.004587, 0.010622) (0.004066, 0.011234)
S6. Motivate Reasoning 0.012354 0.011968 0.012407 0.011929 (0.008901, 0.016671) (0.008133, 0.016052)
S7. Chain-of-thought (COT) 0.013212 0.012862 0.013162 0.012908 (0.009972, 0.017185) (0.009256, 0.016893)
CF Baselines
Discriminative BPR-MF 0.047762 0.060699 0.047742 0.060537 (0.040465, 0.054958) (0.050181, 0.071453)
ItemKNN 0.04905 0.065558 0.048995 0.065367 (0.041122, 0.056398) (0.054682, 0.076526)
NGCF 0.04445 0.053534 0.044376 0.053448 (0.037306, 0.051784) (0.044374, 0.063413)
VAE 0.047158 0.056589 0.047187 0.056502 (0.039967, 0.054677) (0.047093, 0.065803)
LightGCN 0.048295 0.059108 0.048215 0.059019 (0.041621, 0.055491) (0.049809, 0.068851)
TopPop 0.030556 0.032327 0.030539 0.032127 (0.024253, 0.036734) (0.025213, 0.039704)
Additionally, a comparison between the regular performance values (obtained using average across values) and bootstrap averages of model performances provides deeper insights. The bootstrap method, which involves repeatedly resampling with replacement, offers a more robust understanding of the variability in performance measures. Table 5 illustrates this comparison. For instance, while the regular average provides a single estimate of NDCG and Recall, the bootstrap mean extends this by offering confidence intervals. These confidence intervals, represented as Bootstrap Conf. in the table, indicate the range within which the true performance measure is likely to fall with a certain level of confidence.
We outline below the hypotheses regarding personalization that are central to this section:
Hypothesis 1 (H1): The incorporation of a variety of goal-oriented prompts (as detailed in scenarios S1 to S7 in Section 3.1) across diverse settings (e.g., involving different system assignment roles) is anticipated to yield substantially varied performance outcomes in RecLLM. Hypothesis 2 (H2): GPT-based recommenders, particularly in a zero-shot setting, could achieve results comparable to CF baselines in specific scenarios. This is more probable in situations where the system is encouraged to engage in motivated reasoning and/or adopt COT strategies. This hypothesis is founded on observations that such strategies have demonstrated competitive performance in previous ML tasks.
The Impact of Prompt Structure and Setting
In the uppermost tab of Table 5, which presents the settings designated as No Role (R0) for the system role, the results are typically less favorable. Specifically, the prompt scenarios Simple (S1) and Motivated Reasoning (S5) demonstrate better performance within their respective segments. Scenario S1 under No Role (R0) reports NDCG and Recall values of 0.008803 and 0.00994, respectively. In comparison, S5 shows slightly improved metrics, with an NDCG of 0.010152 and a Recall of 0.010906. Overall, these results are the poorest when compared to other settings utilizing LLM-promoted prompts.
Contrastingly, the data in tabs 3 and 4 indicate that when roles are delineated as either Normal Recommender (R1) or Fair Recommender (R2), there is a notable enhancement in both NDCG and Recall metrics. It is of particular note that under these roles (R1 and R2), scenarios such as simple (S1), chain-of-thought (COT) (S7), and Motivated Reasoning (S6) markedly exceed the performance metrics of their No Role (R0) counterparts. The most pronounced performance increase is evident in the fourth tab, where models reach their optimal enhancements. The simple model (S1) advances from 0.008803 to 0.014268 in NDCG (approximately a 62% increase) and from 0.00994 to 0.015796 in Recall (approximately a 59% increase). The Motivated Reasoning (S6) and COT (S7) models also exhibit substantial gains, with S6 escalating from 0.010152 to 0.012354 in NDCG (21.7% increase) and from 0.010906 to 0.011968 in Recall (9.7% increase), and S7 rising from 0.006573 to 0.013212 in NDCG (101% increase) and from 0.005848 to 0.012862 in Recall (120% increase). This trend is similarly observed in the Normal Recommender, Emphasis: Fair (R2/E1) configuration. These results emphasize the significant advantages of integrating system roles into recommendation systems.
Bootstrap values in the R2-E0 setup further affirm the consistency and reliability of these improvements. Both the ‘simple’ and ‘COT’ models exhibit stable performance ranges across different roles and conditions.
Comparison with CF Baselines
When evaluating the effectiveness of GPT-based recommendation models compared to CF baselines, it is apparent that both the NDCG and the recall values for the GPT models are generally lower than those for the CF baselines. This suggests that, while GPT-based models have made significant strides, especially when designated with specific system roles such as “Normal Recommender” (R1) or “Fair Recommender” (R2), they still lag behind CF models in terms of both ranking relevant items at the top (NDCG) and retrieving a broader set of relevant items.
For example, even with the “Fair Recommender” (R2) role applied, the “Simple” (S1) GPT-based model achieves an NDCG of 0.014268 and a recall of 0.015796, which are enhancements of the scenario “No Role” (R0) but still do not surpass the CF baseline models such as “BPR-MF,” “ItemKNN,” “NCF,” and “LightGCN,” that demonstrate higher efficacy in both metrics. The “TopPop” model, while having the lowest scores among the CF baselines, still presents an NDCG and Recall that are competitive with the improved GPT-based models.
In summary, some noteworthy best and worst performances can be found below:
  • For the GPT-based models under the “Fair Recommender” (R2) role with no fairness emphasis (E0):
    • The Simple (S1) scenario achieves the best NDCG of 0.014268 and a Recall of 0.015796.
    • The Diversify (S4) scenario shows the worst NDCG of 0.002844 and a Recall of 0.004901.
  • For the CF baselines:
    • The ItemKNN model has the highest performance with an NDCG of 0.04905 and a Recall of 0.06558.
    • The TopPop model displays the lowest performance among the CF baselines with an NDCG of 0.030556 and a Recall of 0.032327.
Despite the clear impact of prompts on the performance of RecLLM, it can be noted that Top-$k$ recommendation performance is not on par with traditional CF baselines. There are multiple strategies to address this issue, including enhancing prompt clarity and transitioning to few-shot learning. However, we posit that a significant factor contributing to this discrepancy is the possibility that many movies recommended by the GPT-recommender might not be present in the dataset. Therefore, by constraining GPT to select items only from the test item set, we believe that this could markedly improve the accuracy and relevance of the recommendations.
H1: Supported. The introduction of specific roles (e.g., Normal and Fair Recommender) in recommendation systems significantly enhances the performance of generative models. This is evidenced by substantial increases in NDCG and Recall metrics when comparing the No Role scenario (R0) with Role-Based scenarios (R1, R2). It is worth noting in terms of which prompt scenario is effective, the Simple (S1) scenario, suggesting “Recommend 10 movies that the user will likely enjoy,” appears to provide the most effective approach for achieving the best top-$K$ recommendation accuracy. In other scenarios, only the reasoning scenario, such as COT (S7) matches this performance. (H2): Partially Supported. The evidence does not support Hypothesis 2 (H2), which posited that GPT-based recommenders could match CF baselines in certain settings. While GPT models, particularly with system roles and strategies scuh as Simple, ‘Motivated Reasoning’ and ‘COT’, show enhanced NDCG, Recall, they generally do not reach the performance of CF baselines. It is worthwhile that these results indicate a trend towards achieving outcomes closer to those of Top-Pop with the refinement of prompt design, better rationalizing, or transitioning to a few-shot learning setting.
5.2. Stability and Confidence of Personalization Metrics
As stated in Section 3.2, we considered the possibility that external factors such as trending data or algorithmic updates could introduce randomness into responses of GPT-based models. To study the impact of such randomness, each experiment was conducted five times specifically for GPT-based models. It is important to note that prior to our experiments, we lacked specific knowledge about the performance of Chat-GPT prompts. Therefore, our experiments focused exclusively on the R0/E0 scenario, which demonstrated poorer performance compared to the CF baseline. Despite this, the variability of results across different runs is noticeable. Consequently, in the analysis of this section, we should focus on observing the variability of GPT-based models across different runs, rather than comparing them to the CF baselines.
Hypotheses on Stability of the Top-$K$ Accuracy Metric
Hypothesis 1 (H1): Running multiple iterations of GPT-based recommendation queries for a specific user, results in consistent NDCG scores.
The graph in Figures 2 illustrates the NDCG and Recall scores for GPT-based recommendation models across five different runs. These runs overall have been conducted at different points in time that could span from hours to a few days. Notably, there is observable variability in the performance metrics from run to run. For instance,
  • Simple (S1): The NDCG scores are [0.0088, 0.0091, 0.0095, 0.0094, 0.0095], with a std333standard deviation of approximately 0.000305.
  • Motivate reasoning (S6): The NDCG scores are [0.0101, 0.0097, 0.0093, 0.0094, 0.0095], with a std of approximately 0.000316.
  • COT (S7): The NDCG scores are [0.0066, 0.0067, 0.0070, 0.0069, 0.0068], with a std of approximately 0.000158.
Despite the inherent randomness in individual iterations, the general tendency of the results remains consistent, suggesting that while individual runs may produce fluctuating scores, the overall behavior of the models does not significantly deviate. This is evidenced by the low standard deviations in the NDCG scores (0.000305 for S1, 0.000316 for S6, and 0.000158 for S7), indicating a high level of result stability over different runs. Such consistency aligns with the expected behavior of generative models, which can exhibit some degree of randomness in their output due to the stochastic nature of the underlying data/algorithms. However, the close clustering of these scores, as reflected in the narrow stds, indicates a stable pattern of performance.
(a) NDCG@10.
(b) Recall@10
Figure 2. Studying the stability of GPT-based performance metric across different runs.
While the current study demonstrates a high degree of stability in model results over multiple runs, it is crucial to note that these findings do not necessarily guarantee stability over longer time frames, such as weeks or months. The dynamic nature of data and algorithms in generative models means that updates and changes over time could lead to different behaviors and outputs. Nonetheless, caution is required when generalizing these results, and continuous monitoring and periodic re-evaluation of the model’s performance over extended periods are recommended to ensure that the insights remain relevant and accurate.
H1: Supported. The data presented in the NDCG graph aligns with the hypothesis. While there is some expected variability across multiple runs, it remains within an acceptable/predictable range. Confidence intervals overlap significantly for different runs of each GPT-based model. While this consistency across runs confirms the system’s reliability in providing personalized recommendations, it is crucial to note that these findings do not necessarily guarantee stability over longer time frames, such as weeks or months. Therefore, it is advisable to approach these results with caution when considering long-term deployment.
5.3. Diversity and Provider Fairness
This section examines the diversity and fairness of different recommender models. Fairness is assessed by measuring the equitable distribution of recommendations across all items, avoiding bias towards popular or specific subsets (Gini Index, HHI). Diversity is gauged through the variety and uniqueness of the recommendations, ensuring the inclusion of a broad spectrum of items (Entropy), refer to Section 4.2. These evaluations are crucial to understanding each model’s effectiveness in matching user preferences and its capacity to foster a balanced and varied recommendation environment.
H1: Given their training on extensive internet data, GPT-based models are hypothesized to exhibit a greater popularity bias than CF models, potentially leading to lower diversity in recommendations. This will be assessed using the Gini Coefficient and HHI for fairness, and Entropy for diversity. H2: The inclusion of a Fair Recommender as a ”system” role in GPT-based systems is expected to mitigate popularity bias and enhance the diversity of the system’s output.
For H1, upon examining the results presented, in Table 6 a notable observation is that the ItemKNN model consistently exhibits the highest recommendation fairness (achieving the lowest Gini Coefficient and highest Entropy) among all models. This, to conduct a focused discussion, we focused on the remaining machine learning-based CF models and GPT-based models, we will exclude ItemKNN from subsequent comparisons.
Without considering ItemKNN, the general trend suggests that the top-performing GPT models (Simple, Motivate Reasoning, and Chain-of-thought (COT)) are somewhat comparable to the majority of ML-based CF models, except when compared to NGCF, which shows a notably better performance in terms of fairness (Gini Coefficient: 0.950845 vs. 0.0982346, Entropy: 6.420996 vs. 5.042821) and diversity. Intriguingly, methods designed to promote diversification and novelty in GPT case (Surprise and Diversify Recommendation) perform contrarily, evidenced by their increased Gini Coefficients (Surprise: 0.997906, Diversify Recommendation: 0.992349) and reduced Entropy (Surprise: 3.227737, Diversify Recommendation: 4.232139), which is almost the worst in the respective scenarios, suggesting that these strategies may not effectively align with the objectives of reducing bias and enhancing diversity.
Therefore, upon excluding the exceptional performance of ItemKNN, which seems to leverage its distinct mechanism effectively in a dataset characterized by popularity bias, the GPT-based models’ levels of fairness and diversity appear comparable to those of CF models. The data suggests that GPT-based models generally match their CF counterparts, with exceptions such as NGCF which shows superior fairness. Interestingly, NGCF, despite being the least performing CF model in some aspects, still outperforms GPT-based models on fairness. When we turn our attention to the top-performing CF models, namely BPR and LightGCN, it is evident that they achieve their enhanced performance by concentrating user preferences on a narrower selection of items, thus limiting diversity—a trend less likely to be observed in GPT models. This is attributable to the GPT models’ design, which is not inherently influenced by the popularity bias present in our datasets.
For H2, the adaptation of GPT-based models with a ‘Fair Recommender’ system role results in a significant improvement in fairness and diversity metrics. Excluding ItemKNN for its good performance, the top GPT models exhibit better fairness in the Fair Recommender scenario, as evidenced by a noticeable decrease in Gini Coefficient (Simple: from 0.982463 to 0.978925, Chain-of-thought (COT): from 0.986889 to 0.979313) and HHI, coupled with an increase in Entropy (Simple: from 5.042821 to 5.387465, Chain-of-thought (COT): from 4.619500 to 5.365294). This trend suggests that GPT-based models, when guided by the system role within the ChatGPT framework, demonstrate a promising capacity for controllability. Such adaptability could mitigate potential biases inherent in the models—a flexibility that may not be as pronounced in traditional CF models.
However, specific strategies such as Diversification (S4) and Surprise (S5) integrated into GPT-based models do not uniformly follow this positive trend. For example, in the Diversify scenario, we observe a slight increase in the model’s Gini coefficient (from 0.992349 to 0.992603) and a marginal rise in HHI (from 0.034724 to 0.030010), which deviates from the expected reduction. Similarly, the Surprise scenario shows a rise in Gini Coefficient (from 0.997906 to 0.998365) and HHI (from 0.059857 to 0.067952), alongside a reduction in Entropy (from 3.227737 to 3.023948). These anomalies highlight that certain approaches within the Fair Recommender role may inadvertently intensify the biases they aim to alleviate.
The broader implication of this study is that the system role in the ChatGPT framework, alongside certain prompt scenarios, has the potential to align or even surpass the fairness and diversity benchmarks set by classical CF models. In notable instances, the bias is reduced, as indicated by lower Gini and HHI values, and increased Entropy. This indicates a potentially significant step forward in the controllability of GPT-based models to prevent their biases, an attribute that arguably gives them an edge over CF models in managing the diversity and fairness of recommendations.
Answer to H1: Partially supported. The results do not uniformly support H1. While certain GPT-based models do exhibit higher popularity bias as indicated by the Gini Coefficient and HHI compared to some CF models, this is not a consistent finding across all models. For example, models such as NGCF show superior fairness indicators compared to some GPT-based models. However, when excluding the outlier performance of ItemKNN, GPT-based models are generally on par with CF models regarding fairness and diversity. Therefore, the hypothesis that GPT-based models inherently lead to lower diversity due to popularity bias is only partially supported. Answer to H2: Supported. The evidence supports H2 to a significant extent. When GPT-based models adopt a Fair Recommender role, there is a noticeable improvement in fairness and diversity metrics, such as a decrease in Gini Coefficient and HHI, along with an increase in Entropy. This indicates that GPT-based models have an inherent flexibility that, when leveraged with the right system role, can effectively mitigate popularity bias and enhance the diversity of recommendations. In both instances, it is critical to acknowledge that specific strategies, namely “diversification” and “surprise,” do not conform to their anticipated effectiveness—in some cases, they produce outcomes contrary to expectations. This observation implies that certain types of prompts may inadvertently introduce complexity, leading to a confounding of the system’s performance.
Table 6. Diversity and provider fairness of the experimented models
Model Normal Recommender Fair Recommender
Gini Coefficient HHI Entropy Gini Coefficient HHI Entropy
Simple 0.982463 0.017204 5.042821 0.978925 0.010899 5.387465
Genre-focused 0.964743 0.006455 5.919697 0.959879 0.004771 6.110040
Diversify Recommendation 0.992349 0.034724 4.232139 0.992603 0.030010 4.321307
Surprise 0.997906 0.059857 3.227737 0.998365 0.067952 3.023948
Motivate Reasoning 0.981745 0.019189 5.026322 0.979133 0.011218 5.366627
Chain-of-thought (COT) 0.986889 0.027030 4.619500 0.979313 0.011167 5.365294
BPR-MF 0.991758 0.012550 4.658056 0.991758 0.012550 4.658056
Item-KNN 0.914271 0.002877 6.671847 0.914271 0.002877 6.671847
NGCF 0.950845 0.002762 6.420996 0.950845 0.002762 6.420996
VAE 0.989722 0.009554 4.903511 0.989722 0.009554 4.903511
LightGCN 0.989610 0.010546 4.861879 0.989610 0.010546 4.861879
TopPop 0.994859 0.020000 3.912023 0.994859 0.020000 3.912023
5.4. Temporal Freshness Bias in Movie Recommendations
In this section, we study the recency and freshness of the movie recommendations provided by various GPT-based and CF models. The investigation focuses on whether the models demonstrate a bias toward recommending newer or older films. By examining the mean and median release years of the movies suggested by each model, we can gauge their tendency toward recency.
Figure 3. Distribution of Movie Release Years as recommended by different models
Table 7 presents the statistics that inform our understanding of each model’s tendency towards recommending newer or older films. Notably, the GPT-based models ‘Surprise’ and ‘Diversify’ push the boundaries towards recommending more recent items, with median release years of 2006 and 2007. This indicates a clear orientation towards newer movies, possibly interpreting ‘surprise’ and ‘diversification’ in their temporal sense. Since GPT models do not possess inherent information about the novelty of content in terms of user interactions, they likely interpret these directives based on the freshness of the content itself. In contrast, the CF models tend to recommend older films, median release years consistently at 1995, reflecting a stronger inclination towards well-established, historically popular films.
This distinction suggests that GPT-based models might be leveraging their extensive training datasets to prioritize more recent films, which can be beneficial in scenarios where up-to-date recommendations are desired. The broader temporal range of recommendations from GPT-based models, as evidenced by the standard deviation in the year of movies recommended (Surprise: Std Year 10.62, Diversify: Std Year 11.76), further underscores their flexibility in catering to diverse temporal tastes.
Table 7. Model Statistics Reflecting Recency Bias
Model Median Year Std Year
Simple 1999 15.09
Genre-focused 1997 14.99
Diversify 2007 11.76
Surprise 2006 10.62
Motivate reasoning 2002 15.71
COT 1999 15.82
BPR-MF 1995 8.27
ItemKNN 1995 12.14
NGCF 1995 12.25
VAE 1995 8.50
LightGCN 1995 8.34
Pop 1995 5.65
Answer to H1: Contrary to the hypothesis, the results reveals that GPT-based models, especially ‘Diversify’ (Median Year: 2007) and ‘Surprise’ (Median Year: 2006), tend to recommend significantly newer movies than CF models. CF models such as BPR-MF (Median Year: 1995) and Pop (Median Year: 1995) generally suggest older films, as reflected in the lower mean and median years. The larger standard deviation in years for GPT-based models like ‘Surprise’ (Std Year: 10.62) also indicates a broader temporal range in their recommendations. This suggests that GPT-based models, perhaps due to their ability to capture more recent trends and discussions from diverse datasets, are capable of recommending more recent content than CF models, thus refuting the initial hypothesis.
5.5. Analysis of Genre Bias/Affinity in Recommender Systems
The goal of this section is to examine the genre biases present in movie recommendations provided by Collaborative Filtering (CF) models and GPT-based models. By comparing the prevalence of genre terms in the recommendations, we aim to understand if a particular model, or class of models, has a specific genre affinity (or bias) which could illuminate the performance differences between these models.
H1: We hypothesize that the differences in recommendation performance between CF and GPT-based models may be attributed to their respective genre preferences. CF models are expected to closely adhere to genre affinity in users’ profiles, while GPT models may exhibit a predominant recommendation of certain genres over others.
Upon analyzing the word clouds, it is evident that CF models such as BPR-MF, LightGCN, and VAE predominantly recommend genres such as “Action,” “Adventure,” “Thriller,” and “Sci-Fi.” These genres appear frequently and in larger font sizes, indicating their dominance in the recommendations made by these models.
In contrast, GPT-based models, represented in the lower word clouds, show a more varied genre distribution with “Comedy,” “Drama,” and “Romance” and even “Thriller” and “Adventure” being more prominent. These results are interesting and suggest that GPT-based RecLLMs may be less constrained by the users’ historical genre affinities and more exploratory in their recommendations, possibly leading to a broader but potentially less precise genre match to individual user profiles.
Figure 4. WordCloud of Movie Genres as recommended by different models. Top model correspond to CF models (BPR-MF, LightGCN, RecVAE), while lower models include GPT-based recommenders (Simple, Diversity, COT).)
Answer to H1: The analysis supports the hypothesis to a certain extent. While CF models tend to reinforce existing user preferences by frequently recommending popular genres such as “Action” and “Sci-Fi,” GPT-based models demonstrate a broader genre coverage in their recommendations. The presence of a more diverse set of genres like “Comedy,” “Drama,” and “Romance” in GPT-based recommendations suggests these models may be utilizing their extensive training on diverse data sources to introduce a wider variety of content. This could potentially lead to a discovery of new interests for users, though it may sometimes diverge from their established preferences. Such a tendency for broader genre representation by GPT-based models could be a contributing factor to their distinct performance characteristics compared to CF models.
5.6. Implicit vs. explicit scenario
Analyzing the impact of including ratings in the prompt shows noteworthy results. In our study, we initially avoided using ratings for the GPT models, aligning with the implicit setting of the Collaborative Filtering (CF) baseline. However, for completeness, we explored the effect of revealing user ratings for selected items.
The “Include-rating” model demonstrates a significant improvement compared to other GPT-based scenarios. With an NDCG of 0.013230 and a Recall of 0.014721, it surpasses genre-focused, Diversify Recommendation, Surprise, Motivate Reasoning, and chain-of-thought (COT) models. The Bootstrap NDCG and Recall for the Include-rating model are 0.013280 and 0.014673, respectively, indicating consistency in performance. The confidence intervals for NDCG (0.009737 to 0.017012) and Recall (0.010261 to 0.019795) further emphasize its reliability.
These results suggest that including ratings significantly enhances the recommendation system’s performance. It provides a more personalized and accurate reflection of user preferences, which is a key aspect in recommendation systems. While other GPT-based models showed varied performance, the inclusion of ratings seems to offer a more consistent and improved outcome. This insight could be valuable in refining other scenarios and models, potentially leading to better overall system performance.
6. Conclusion and Future Directions
This research presents a comprehensive analysis of Recommender Systems using Large Language Models (RecLLMs), specifically focusing on ChatGPT-based systems. The study emphasizes the nuanced capabilities and inherent biases of these systems, highlighting their contrasting behaviors compared to traditional CF in the domain of movie recommendations. Our experimental findings underscore the significant impact of ‘prompt design’ strategies on recommendation quality, touching upon aspects such as accuracy, provider fairness, diversity, temporal stability, genre dominance, and freshness (recency).
We discovered that role-based prompts (those using clear system roles such as you a Recommender, or you are a Fair recommender) enhance fairness and diversity, effectively mitigating popularity bias. Interestingly, GPT-based models, while not always paralleling the performance of CF baselines, exhibit a preference for recommending newer and more diverse movie genres.
Moreover, our experiments reveal that including ratings in prompts appears to improve the personalization aspect of recommendations. GPT-based models also have a distinct inclination towards recommending newer, particularly those released post-2000. We hypothesize that a key reason for certain discrepancies in recommendations quality is that the GPT model might suggest movies not included in the dataset. By limiting the GPT model to choose only from a predefined set of test items, we anticipate a notable improvement in both the accuracy and relevance of its recommendations. Additionally, transitioning to a zero-shot scenario could significantly enhance personalization results in this context. In genral, we think that it important to balance personalization power of CF models with the broader exploratory nature of GPT-based RecLLMs, to ensure a diverse and engaging user experience.
Key areas for future research include:
  1. (1)
    Refine Prompt Design Strategies and Dataset Constraints. This involves developing prompts that effectively balance user preferences with broader content exploration, while also constraining the system to select items only from within the dataset. This limitation is expected to improve the accuracy and relevance of recommendations.
  2. (2)
    Explore Few-shot Learning to Improve accuracy and Contextual Relevance. Few-shot learning involves exposing the model to a small number of examples from the target domain, which can enhance its ability to make more informed recommendations.
  3. (3)
    Integrate GPT with CF in a RAG framework. Aimed to create a hybrid approach that leverages the strengths of both methodologies. This combination could potentially lead to more robust and effective recommender systems.
  4. (4)
    Expand research to cross-domain applications and conduct long-term user engagement studies. This includes evaluating RecLLMs in various content domains (e.g., books, music, e-commerce) and studying user interaction over extended periods to gauge satisfaction, trust, and the evolving nature of user preferences.
By focusing on these key areas, future research can significantly advance the field of recommender systems, particularly in leveraging the capabilities of large language models like GPT in diverse and dynamic recommendation scenarios.